OpenStudy (sparkk):

One last question :) I will medal and fan :) Thank you for helping if you do! Solve for x: −3|x + 7| = −12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (sparkk):

I say no solution?

OpenStudy (sparkk):

???

OpenStudy (holsteremission):

For any number \(z\), \(|z|\) is always positive. Your equation has a negative number multiplied by a positive number (the left side), which is always going to give a negative number. This guarantees that there's at least one solution.

OpenStudy (holsteremission):

To find the solution(s), you can write \[\begin{align*} -3|x+7|&=-12\\[1ex] |x+7|&=4 \end{align*}\]Then, invoking the definition of the absolute value function, you can analyze the left side a bit more: \[|x|=\begin{cases}x&\text{if }x\ge0\\[1ex]-x&\text{if }x<0\end{cases}\implies|x+7|=\begin{cases}x+7&\text{if }x+7\ge0\\[1ex]-(x+7)&\text{if }x+7<0\end{cases}=\begin{cases}x+7&\text{if }x\ge-7\\[1ex]-x-7&\text{if }x<-7\end{cases}\]Consider one case at a time. (1) If \(x\ge-7\), then the equation becomes \[|x+7|=4\implies x+7=4\] (2) If \(x<-7\), then you get \[|x+7|=4\implies -x-7=4\]and you can see that there are indeed exactly two (not zero) solutions.

OpenStudy (holsteremission):

Oops, copying the cases that are cut off: \[\cdots\implies|x+7|=\begin{cases}x+7&\text{if }x+7\ge0\\[1ex]-(x+7)&\text{if }x+7<0\end{cases}=\begin{cases}x+7&\text{if }x\ge-7\\[1ex]-x-7&\text{if }x<-7\end{cases}\]